What is the equation parallel to the line 3x+2y=6

Answer by TimothyLamb(4379)

Question:
What is the equation parallel to the line 3x+2y=6
---
there are an infinite number of lines parallel to the above line, not just one parallel line as the question implies with this statement: "the line".
---
convert the equation to slope-intercept form:
---
3x + 2y = 6
2y = -3x + 6
y = (-3/2)x + 3
---
the set of parallel lines have this equation : y = (-3/2)x + b
they all have the same slope m = -3/2
but different y-intercepts, for example b= 0, b= 1, b= -2764.39757, etc.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php